OpenStudy (anonymous):

By what number must you multiply each side of the equation 4/5 x =8 to produce the equivalent equation x = 10?

OpenStudy (anonymous):

I think it is 4/5 that you have to multiply to both sides but I need to know

OpenStudy (anonymous):

You need to combine the equations 4/5 x = 8 and x = 10 (or 5/5x = 10) This leaves you with 9/5 x = 18. Now solve for x.

OpenStudy (noorik.):

you need to multiply by 5/4 because 5/4 * 4/5 = 1